# ActionPiece: Contextually Tokenizing Action Sequences for Generative Recommendation

### v1

@@ B.1 Vocabulary Construction Algorithm
12:        end for# Process all ordered token pairs between $A_{k}$ and $A_{k+1}$
13:        if $k<t-1$ then
14:           $\mathcal{A}_{k+1}\leftarrow S^{\prime}[k+1]$
15:           for all $c_{i}\in\mathcal{A}_{k},c_{j}\in\mathcal{A}_{k+1}$ do
16:              $\text{count}(c_{i},c_{j})\leftarrow\text{count}(c_{i},c_{j})+1/(|\mathcal{A}_{k}|\times|\mathcal{A}_{k+1}|)$ weight of tokens from two adjacent sets (Equation 2)
17:           end for
18:        end if
19:     end for
@@ C.2 Time Complexity
Appendix D Discussion: Comparison Between ActionPiece and BPE
Table 5: Comparison between ActionPiece and BPE.
Aspect BPE ActionPiece
Data Type text sequences action (unordered feature set) sequences
Token a byte sequence a feature set
Initial Vocabulary single bytes single-feature sets
While ActionPiece follows a similar algorithmic framework as BPE, its design is fundamentally different because it is tailored for tokenizing action sequences. To clarify, we summarize the key differences in Table 5.
Appendix E Datasets
Datasets #Users #Items #Actions Avg. $t$
Sports 18,357 35,598 260,739 8.32
Beauty 22,363 12,101 176,139 8.87
CDs 75,258 64,443 1,022,334 14.58
@@ Appendix E Datasets
Categories. Among all the datasets, “Sports” and “Beauty” are two widely used benchmarks for evaluating generative recommendation models (Rajput et al., 2023; Jin et al., 2024; Hua et al., 2023). We conduct experiments on these benchmarks to ensure fair comparisons with existing results. Additionally, we introduce “CDs”, which contains about 4×4$\times$4 × more interactions than “Sports”, making it a larger dataset for evaluating the scalability of GR models. For “CDs”, we apply the same data processing strategy as the public benchmarks. The statistics of the processed datasets are shown in Table 6.
@@
Item text features. Following Rajput et al. (2023); Zheng et al. (2024a); Sheng et al. (2025), the first step for feature engineering is to combine multiple raw text features into a single sentence for each item. Then, we use a pretrained sentence embedding model to encode this sentence into a vector representation. In all our implementations, we concatenate title, price, brand, feature, categories, and description, and use sentence-t5-base (Ni et al., 2022) as the sentence embedding model.
@@
  • We quantize the sentence embeddings using residual quantization (RQ) (Rajput et al., 2023; Zeghidour et al., 2021; Zheng et al., 2024b) into three codes, each with $256$ candidates. To prevent conflicts, we add an extra identification code. These four codes together serve as the RQ-based semantic IDs for TIGER and SPM-SID.
@@ Appendix F Baselines
F.1 ID-Based Sequential Recommendation Methods
@@
F.2 Feature-Enhanced Sequential Recommendation Methods
@@
F.3 Generative Recommendation Methods
@@
Appendix G Implementation Details
@@
Appendix H Reproduction
To improve reproducibility, we provide the algorithms of vocabulary construction and segmentation processes in Algorithms 1, 2, 3 and 4. We also provide the pseudocode for the efficient vocabulary construction implementation in Figure 7. In addition, we provide the best hyperparameters of ActionPiece for all experimental datasets in Table 7.

### v2

@@ B.1 Vocabulary Construction Algorithm
12:        end for
13:        # Process all ordered token pairs between $A_{k}$ and $A_{k+1}$
14:        if $k<t-1$ then
15:           $\mathcal{A}_{k+1}\leftarrow S^{\prime}[k+1]$
16:           for all $c_{i}\in\mathcal{A}_{k},c_{j}\in\mathcal{A}_{k+1}$ do
17:              $\text{count}(c_{i},c_{j})\leftarrow\text{count}(c_{i},c_{j})+1/(|\mathcal{A}_{k}|\times|\mathcal{A}_{k+1}|)$ Weight of tokens from two adjacent sets (Equation 2)
18:           end for
19:        end if
@@ C.2 Time Complexity
Appendix D Additional Related Work
Aligning LLMs with user actions. A key motivation for developing action tokenization methods is to provide an efficient and effective way of aligning pretrained generative models (e.g., LLMs (OpenAI, 2022; Anil et al., 2023; Touvron et al., 2023; Zhao et al., 2023)) with user action data (Geng et al., 2022; Zheng et al., 2024; Hou et al., 2025). A typical pipeline involves tokenizing action sequences into action tokens and then performing instruction tuning on inputs that combine both language and action tokens. The choice of tokenization strategy plays a critical role. We identify three main paradigms:
  • One token per item: Each action is represented by a single token, resulting in a dense vector that is typically derived from a pretrained semantic encoder or a learnable embedding table (Hou et al., 2022; Liao et al., 2024; Zhu et al., 2024b; Zhang et al., 2025b; Kim et al., 2024). While this approach is efficient in terms of sequence length, it suffers from memory and scalability issues, particularly since the number of unique items often exceeds the typical vocabulary size of LLMs. Aligning LLMs with such large vocabularies introduces both engineering and optimization challenges.
  • Text-based tokenization: Each action is expressed as a textual string, which naturally aligns with the LLMs’ native modality (Geng et al., 2022; Zhang et al., 2025a; Bao et al., 2023). However, this leads to substantially longer token sequences, resulting in inefficiencies during tokenization and increased inference latency.
  • Discrete tokens: Actions are tokenized into a small number of discrete tokens drawn from a compact, shared vocabulary (Zheng et al., 2024; Tan et al., 2024; Jin et al., 2024). This approach balances sequence length and memory efficiency, making it a practical choice for building LLM-based recommender systems.
Appendix E Additional Discussions
E.1 Benefits of Set Permutation Regularization
SPR benefits the model from multiple perspectives:
  • Token utilization perspective. SPR effectively prevents the features of a single action from being consistently merged into the most compressed (high-level) tokens. Instead, it allows the action to be tokenized into both high-level and low-level tokens, depending on the permutation and token merge rules. This increases the number of tokens actively involved during both training and inference. As shown in Figure 5 and discussed in Section 4.4.2, SPR significantly improves token utilization - from 56.89% to 95.33% by the 5th epoch - indicating that a greater proportion of tokens are trained after applying SPR.
  • Data augmentation perspective. From the perspective of data augmentation, SPR enriches the token sequences available for model training. Without SPR, each action sequence can only be tokenized into a single, fixed token sequence. In contrast, SPR allows each action sequence to be tokenized in multiple ways (as shown in Figure 1). While these augmented sequences preserve the same semantic information, they expose the model to richer token patterns. Training on these diverse token sequences helps the model generalize better, as evidenced by the performance of variant (3.1) in Table 5.
  • Ensemble perspective. SPR also enables inference-time augmentation. A given input action sequence can be augmented into multiple token sequences during inference. Each sequence may yield a different ranking of the next possible items. By ensembling these recommendation results, overall performance can be enhanced, as demonstrated by variant (3.2) in Table 5 and further illustrated in Figure 6.
E.2 Connections Between ActionPiece and Feature Crossing
A core component of ranking models is automatically capturing feature-crossing patterns that are helpful for recommendation performance (Shan et al., 2016; Wang et al., 2017, 2021; Bian et al., 2022), which is conceptually related to ActionPiece. The key difference lies in the level at which interactions are modeled. Prior works typically perform automatic feature crossing at the model level, learning interactions implicitly through the network architecture. In contrast, ActionPiece merges features at the data level.
Although we do not directly compare against these specific feature-crossing models, we include several baselines that follow a similar design philosophy. For example, HSTU in Table 4 and variant (2.1) in Table 5 use the same underlying item features as ActionPiece but input them in a flattened form, without merging. In these setups, feature interactions are expected to be learned by the autoregressive model via self-attention and feed-forward layers, i.e., at the model level.
Our results demonstrate that ActionPiece, which performs data-level feature merging, consistently outperforms these methods in both recommendation quality (Tables 4 and 5) and efficiency (Figure 4). This is particularly evident in normalized sequence length (NSL): both HSTU and variant (2.1) yield NSL values of 1, reflecting longer sequences compared to the more compact sequences tokenized by ActionPiece.
Appendix F Datasets
@@
Categories. Among all the datasets, “Sports” and “Beauty” are two widely used benchmarks for evaluating generative recommendation models (Rajput et al., 2023; Jin et al., 2024; Hua et al., 2023). We conduct experiments on these benchmarks to ensure fair comparisons with existing results. Additionally, we introduce “CDs”, which contains about 4×4$\times$4 × more interactions than “Sports”, making it a larger dataset for evaluating the scalability of GR models. For “CDs”, since there are no publicly available results from generative recommendation methods to date, we closely followed the experimental settings used in public benchmarks like “Sports” and “Beauty” to ensure fair comparisons.
@@
Item text features. Following Rajput et al. (2023); Zheng et al. (2024); Hou et al. (2024a); Sheng et al. (2025), the first step for feature engineering is to combine multiple raw text features into a single sentence for each item. Then, we use a pretrained sentence embedding model to encode this sentence into a vector representation. In all our implementations, we concatenate title, price, brand, feature, categories, and description, and use sentence-t5-base (Ni et al., 2022) as the sentence embedding model.
@@
  • We quantize the sentence embeddings using residual quantization (RQ) (Rajput et al., 2023; Zeghidour et al., 2021; Zheng et al., 2025) into three codes, each with $256$ candidates. To prevent conflicts, we add an extra identification code. These four codes together serve as the RQ-based semantic IDs for TIGER and SPM-SID.
@@ Appendix G Baselines
G.1 ID-Based Sequential Recommendation Methods
@@
G.2 Feature-Enhanced Sequential Recommendation Methods
@@
G.3 Generative Recommendation Methods
@@
Appendix H Implementation Details
@@
Inference. The inference process of ActionPiece follows the same procedure as TIGER. The decoder autoregressively generates token sequences representing the target items. During training, we use the original item features as labels, without any augmentation or token merging. At inference time, for each augmented version of a test case, we apply beam search to generate top-ranked token sequences. The most probable sequences (i.e., prefixes) are retained in the beam (with beam size specified in Table 7), and the model continues generating tokens one at a time until the target sequence length is reached.